# Functional Determinants for Constrained Path Integrals in Minisuperspace Jackiw-Teitelboim Gravity

### v1

@@ 6.1 Bianchi IX Path Integral and Constraint Structure
which coincides with the on-shell action obtained in Ref. DiazDorronsoro:2018wro.
@@ 6.2 Functional Determinant
which agrees with the propagator obtained in Ref. DiazDorronsoro:2018wro.  [^3] Our convention keeps the overall factor $2\pi^{2}$ in the action (61), whereas Ref. DiazDorronsoro:2018wro absorbs this factor into its normalization conventions. Therefore, by shifting $\hbar\to 2\pi^{2}\hbar$, we obtain an expression that matches exactly.  [^4] According to the author’s comment, the prefactor in Ref. DiazDorronsoro:2018wro was given by the Van Vleck determinant formula.
In summary, evaluating the functional determinant via the Gelfand-Yaglom theorem fixes the prefactor and thereby validates our treatment of the constrained path integral. Finally, integrating the fixed-lapse propagator (94) over the positive lapse yields the full Lorentzian transition amplitude. A detailed analysis of the associated thimble structure for the Bianchi IX lapse integral, including several different boundary conditions, can be found in Refs. DiazDorronsoro:2018wro; Feldbrugge:2018gin; Janssen:2019sex; Lehners:2024kus, so we will not discuss it further here.
@@ 7 Conclusion
For Jackiw-Teitelboim gravity, the constrained formulation reproduces the propagator previously obtained by performing the dilaton and scale-factor path integrals explicitly Honda:2024hdr, providing a nontrivial consistency check of the delta-functional treatment and the associated prefactor. Upon integrating over the lapse, we derived closed-form expressions in terms of Bessel/Hankel functions and reviewed the resulting saddle-point structure Honda:2024hdr, which separates an oscillatory, semiclassical regime from an exponentially suppressed, genuinely quantum regime. We also extended the analysis to a quadratic dilaton potential $U(\phi)=m^{2}\phi^{2}$. In this case, the determinant prefactor is unchanged, and the qualitative thimble structure is preserved.
As an application beyond Jackiw-Teitelboim gravity, we applied the same strategy to biaxial Bianchi IX minisuperspace. Integrating out one minisuperspace variable produces a functional constraint for the remaining scale factor degree of freedom. We evaluated the associated functional determinant via the Gelfand-Yaglom method and obtained the fixed-lapse propagator including its prefactor. Our result agrees with the propagator reported in Ref. DiazDorronsoro:2018wro.
@@ Acknowledgments
H. M. would like to thank Masazumi Honda, Kota Numajiri, Oliver Janssen, Kazumasa Okabayashi, Daiki Saito and Takahiro Terada for useful discussions regarding the minisuperspace analysis of Jackiw-Teitelboim gravity and Bianchi IX quantum cosmology. This work is supported by JSPS KAKENHI Grant No. JP23K13100.

### v2

@@ 6.1 Bianchi IX Path Integral and Constraint Structure
which coincides with the on-shell action obtained in Ref. Diaz Dorronsoro et al. (2018).
@@ 6.2 Functional Determinant
which agrees with the propagator obtained in Ref. Diaz Dorronsoro et al. (2018).  [^4] Our convention keeps the overall factor $2\pi^{2}$ in the action (61), whereas Ref. Diaz Dorronsoro et al. (2018) absorbs this factor into its normalization conventions. Therefore, by shifting $\hbar\to 2\pi^{2}\hbar$, we obtain an expression that matches exactly.  [^5] According to the author’s comment, the prefactor in Ref. Diaz Dorronsoro et al. (2018) was given by the Van Vleck determinant formula.
In summary, evaluating the functional determinant via the Gelfand-Yaglom theorem fixes the prefactor and thereby validates our treatment of the constrained path integral. Finally, integrating the fixed-lapse propagator (94) over the positive lapse yields the full Lorentzian transition amplitude. A detailed analysis of the associated thimble structure for the Bianchi IX lapse integral, including several different boundary conditions, can be found in Refs. Diaz Dorronsoro et al. (2018); Feldbrugge et al. (2018); Janssen et al. (2019); Lehners (2025), so we will not discuss it further here.
@@ 7 Conclusion
For Jackiw-Teitelboim gravity, the constrained formulation reproduces the propagator previously obtained by performing the dilaton and scale-factor path integrals explicitly Honda et al. (2025), providing a nontrivial consistency check of the delta-functional treatment and the associated prefactor. Upon integrating over the lapse, we derived closed-form expressions in terms of Bessel/Hankel functions and reviewed the resulting saddle-point structure Honda et al. (2025), which separates an oscillatory, semiclassical regime from an exponentially suppressed, genuinely quantum regime. We also extended the analysis to a quadratic dilaton potential $U(\phi)=m^{2}\phi^{2}$. In this case, the determinant prefactor is unchanged, and the qualitative thimble structure is preserved.
As an application beyond Jackiw-Teitelboim gravity, we applied the same strategy to biaxial Bianchi IX minisuperspace. Integrating out one minisuperspace variable produces a functional constraint for the remaining scale factor degree of freedom. We evaluated the associated functional determinant via the Gelfand-Yaglom method and obtained the fixed-lapse propagator including its prefactor. Our result agrees with the propagator reported in Ref. Diaz Dorronsoro et al. (2018).
@@ Acknowledgments
H. M. would like to thank Masazumi Honda, Oliver Janssen, Kota Numajiri, Kazumasa Okabayashi, Daiki Saito, and Takahiro Terada for useful discussions regarding the minisuperspace analysis of Jackiw-Teitelboim gravity and Bianchi IX quantum cosmology. This work is supported by JSPS KAKENHI Grant No. JP23K13100.